# Digital Twin Calibration for Biological System-of-Systems: Cell Culture Manufacturing Process

### v1

@@ 3 Digital Twin Calibration for Bio-SoS Mechanistic Model
To facilitate the mechanistic model digital twin calibration, in this paper, we consider the physical system as a finite horizon stochastic process. Its dynamics can be characterized by a Bio-SoS mechanistic model specified by Eq. (2) and (3) with underlying true parameters, denoted by $\boldsymbol{\theta}^{c}=[\boldsymbol{\alpha}_{0}^{c},\boldsymbol{\alpha}_{1}^{c},\boldsymbol{\beta}^{c}]$. Our goal is to develop an optimal learning calibration approach that can efficiently guide the sequential design of experiments and most informative data collection to improve model fidelity and the prediction accuracy of the digital twin. For simplification, we consider batch-based cell culture experiments with the selection of initial concentration $\boldsymbol{s}_{0}$ as the decision variable.
The trajectory $\tau=(\boldsymbol{s}_{0},\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{T})$ over a horizon of $T$ time steps has the joint probability density function
$p(\boldsymbol{\tau}|\boldsymbol{\theta},\boldsymbol{s}_{0})=\prod_{t=0}^{T-1}p(\boldsymbol{s}_{t+1}|\boldsymbol{s}_{t};\boldsymbol{\theta})\quad\text{ with $\boldsymbol{s}_{0}=\pi(\boldsymbol{\omega})$,}$ (6)
where $\pi(\boldsymbol{\omega})$ is a deterministic design or calibration policy parameterized by $\boldsymbol{\omega}$. Each transition probability $p(\boldsymbol{s}_{t+1}|\boldsymbol{s}_{t};\boldsymbol{\theta})$ follows a normal distribution as defined in Eq. (5).
At each $k$-th calibration iteration, we select one design following the latest policy, i.e. $\boldsymbol{s}_{0}^{(k)}=\pi(\boldsymbol{\omega}_{k})$, run an experiment to collect one sample path from the physical system, and then update the parameter $\widehat{\boldsymbol{\theta}}_{k}$. Specifically, given the historical samples $\mathcal{D}_{k}=\{\boldsymbol{\tau}_{n}\}_{n=1}^{k}$, we consider maximizing the log-likelihood for model parameters $\boldsymbol{\theta}$, i.e.,
@@
For digital twin calibration, we select the design of new experiment at the beginning of $k$-th iteration. This selection is based on the evaluation of our current model on a pre-specified set of evaluation and prediction points $\mathcal{D}_{test}=\{\boldsymbol{s}_{0}^{h}\}_{h=1}^{H}$. The objective is to minimize the MSE of digital twin prediction,
@@
where $\boldsymbol{u}_{k}(\widehat{\boldsymbol{\theta}}_{k-1},\boldsymbol{\omega})$ denote the virtual model update function with the simulated trajectory $\boldsymbol{\tau}^{\text{d}}_{k}$ (see Eq. (10)) and $Y=\phi(\boldsymbol{s}_{T})$ represents the system outputs of interest, such as yield and product quality attributes. The superscript "p" indicates the output from the physical system and "d" for digital twin.
To sample efficiently guide digital twin calibration, we propose a gradient-based optimal learning approach with the procedure illustration as shown in Figure 2 and develop a calibration algorithm summarized in Algorithm 1. In specific, for the optimization problem (7) , we use the gradient-based method to search for the MLE, i.e.,
$\widehat{\boldsymbol{\theta}}_{k}\leftarrow\widehat{\boldsymbol{\theta}}_{k-1}+\text{Adam}_{k}\left(\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{\tau}_{k};{\boldsymbol{\theta}},\boldsymbol{s}_{0}^{(k)})\mid_{\boldsymbol{\theta}=\widehat{\boldsymbol{\theta}}_{k-1}}\right),$ (9)
@@
$\boldsymbol{u}_{k}(\widehat{\boldsymbol{\theta}}_{k-1},\boldsymbol{\omega})=\widehat{\boldsymbol{\theta}}_{k-1}+\text{Adam}_{k}\left(\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{\tau}_{k}^{\text{d}};\boldsymbol{\theta},\boldsymbol{\omega})\mid_{\boldsymbol{\theta}=\widehat{\boldsymbol{\theta}}_{k-1}}\right)$ (10)
@@
Figure 2: The procedure illustration of the proposed digital twin calibration approach.
@@
Since $\boldsymbol{\theta}^{c}$ is unknown, we use the bootstrapping to quantify the model parameter estimation uncertainty in Section 5. Following the upper part of Figure 2, the prediction error propagates from $\boldsymbol{\omega}$ to model parameter update $\boldsymbol{u_{k}}$, and then to our objective $J_{k}(\boldsymbol{\omega})$. By taking the gradient of $J_{k}(\boldsymbol{\omega})$, we can discern the influence of the experimental design specified by $\boldsymbol{\omega}$ on the final prediction error, which assists in selecting a more effective design. Since deriving an exact expression for this gradient is complex, we develop a linear noise approximation closed-form representation, which will be detailed in Section 5. This approximation provides an efficient method for improving calibration policy.
For $k=1:K$
  1. 1.
  • 2.
  • Algorithm 1 Gradient-based Optimal Learning
@@ 4 Model Inference and Parameter Update
    The gradient for the first term is $\nabla_{\boldsymbol{\theta}}\log|\Sigma(\boldsymbol{\theta})|=\text{tr}\left(\Sigma(\boldsymbol{\theta})^{-1}\nabla_{\boldsymbol{\theta}}\Sigma(\boldsymbol{\theta})\right),$ where $\nabla_{\boldsymbol{\theta}}\Sigma(\boldsymbol{\theta})$ is a tensor of $M\cdot M\cdot q$ with $q$ as the length of $\boldsymbol{\theta}$, and "tr" represents the trace for first two dimensions. The gradient for the second term is
    ${\nabla_{\boldsymbol{\theta}}}(\boldsymbol{s}_{t+1}-\boldsymbol{s}_{t}-\boldsymbol{\mu}(\boldsymbol{\theta}))^{\top}\Sigma(\boldsymbol{\theta})^{-1}(\boldsymbol{s}_{t+1}-\boldsymbol{s}_{t}-\boldsymbol{\mu}(\boldsymbol{\theta}))$
    $=$ $-2\nabla_{\boldsymbol{\theta}}{\boldsymbol{\mu}}(\boldsymbol{\theta})^{\top}\Sigma(\boldsymbol{\theta})^{-1}(\boldsymbol{s}_{t+1}-\boldsymbol{s}_{t}-\boldsymbol{\mu}(\boldsymbol{\theta}))+(\boldsymbol{s}_{t+1}-\boldsymbol{s}_{t}-\boldsymbol{\mu}(\boldsymbol{\theta}))^{\top}\left(-\Sigma(\boldsymbol{\theta})^{-1}{\nabla_{\boldsymbol{\theta}}\Sigma(\boldsymbol{\theta})}\Sigma(\boldsymbol{\theta})^{-1}\right)(\boldsymbol{s}_{t+1}-\boldsymbol{s}_{t}-\boldsymbol{\mu}(\boldsymbol{\theta})).$
    Given the calculated gradient, we can use the Adam optimizer Kingma and Ba (2014) for updating $\widehat{\boldsymbol{\theta}}$. Following the algorithm in their paper, for the $k$-th iteration, we have
@@
    where $\lambda$ is the learning rate, $G_{l}=\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{\tau}_{l};\boldsymbol{\theta},\boldsymbol{s}_{0}^{(l)})\mid_{\boldsymbol{\theta}=\widehat{\boldsymbol{\theta}}_{l-1}}$ is the gradient of the likelihood function at $l$-th iteration, $g_{l}^{2}$ is the element-wise square, $\xi_{1}=0.9,\xi_{2}=0.999$ are hyper-parameters related to exponential decay rates, $\epsilon=10^{-8}$ is a small number to avoid zero in the denominator.
@@ 5 Policy Gradient Estimation and Optimal Learning
    In this section, we turn our focus on learning the calibration policy by minimizing the prediction MSE in (8). To develop a surrogate model of this MSE, we employ the linear noise approximation in Section 5.1 to approximate the stochastic dynamics of the system and the solution of SDE-based mechanistic model, reflecting its sensitivity to initial conditions and parameter values. This approach allows us to describe the system state at any given time as a normally distributed random variable, facilitating the derivation of analytical expressions for the gradient estimator. To keep the mechanism information, in Section 5.2, a first-order Euler’s Method is utilized to derive the closed-form solutions of the SDEs based mechanisms obtained from linear noise approximation.
@@ 5.1 Linear Noise Approximation on Bio-SoS Dynamics
    We use linear noise approximation (LNA) (Fearnhead et al., 2014) to derive the solution of SDEs as shown in Lemma 2 that allows us to analyze the estimation error propagation from ${\boldsymbol{\theta}}$ to the output prediction and develop a surrogate model of the MSE (8). Conditional on $\boldsymbol{s}_{0}$ and $\boldsymbol{\theta}$, we denote the state of systems at time $t$ as $\boldsymbol{s}_{t}(\boldsymbol{\theta})$. Considering the SDEs characterizing the Bio-SoS dynamics as shown in (4), we have
@@
    Obtaining an analytical solution for the series of ODEs described in (15) is challenging. Therefore, we use a numerical solution approach, as described in Section 5.2. This method involves applying the solution iteratively at each time step $t$, allowing us to approximate the system’s behavior over time. For the physical system and digital twin, we denote
    $\boldsymbol{s}_{t}(\boldsymbol{\theta}^{c})\sim N\left(\boldsymbol{\eta}_{t},\boldsymbol{\Psi}_{t}\right),\quad\boldsymbol{s}_{t}({\boldsymbol{u}}_{k})\sim N\left(\widehat{\boldsymbol{\eta}}_{t},\widehat{\boldsymbol{\Psi}}_{t}\right).$
    Then we have $\boldsymbol{s}_{T}(\boldsymbol{\theta}^{c})-\mathbf{s}_{T}({\boldsymbol{u}}_{k})\sim N\left(\boldsymbol{\eta}_{T}-\widehat{\boldsymbol{\eta}}_{T},\boldsymbol{\Psi}_{T}+\widehat{\boldsymbol{\Psi}}_{T}\right).$ For any $m$-th component $Y$, we have a surrogate model of the MSE (8)
    ${\mathbb{E}}\left[\left(Y^{\text{p}}(\boldsymbol{\theta}^{c})-Y^{\text{d}}({\boldsymbol{u}}_{k})\right)^{2}\right]$ $=\left[{\mathbb{E}}\left(Y^{\text{p}}(\boldsymbol{\theta}^{c})-Y^{\text{d}}({\boldsymbol{u}}_{k})\right)\right]^{2}+{\rm Var}\left(Y^{\text{p}}(\boldsymbol{\theta}^{c})-Y^{\text{d}}({\boldsymbol{u}}_{k})\right)$
    $=\left[\boldsymbol{\eta}_{T,m}-\widehat{\boldsymbol{\eta}}_{T,m}\right]^{2}+\boldsymbol{\Psi}_{T,mm}+\widehat{\boldsymbol{\Psi}}_{T,mm}.$ (16)
@@ 5.2 First-Order Euler’s Method for Solving LNA
    Then conditional on $\widehat{\boldsymbol{\theta}}_{k-1}$, by applying a chain rule, we can update the policy parameters by
@@
    where $\tilde{G}_{k}=\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{\tau}_{k}^{\text{d}};\boldsymbol{\theta},\boldsymbol{s}_{0}^{(k)})\mid_{\boldsymbol{\theta}=\widehat{\boldsymbol{\theta}}_{k-1}}$ and $\tilde{G}_{t}=G_{t},\text{for }t=1,\ldots,k-1$. Here $\tilde{G}_{k}$ is a gradient based on simulation trajectory $\boldsymbol{\tau}_{k}^{\text{d}}$ since $\boldsymbol{\omega}_{k}$ is updated at the beginning of $k$-th iteration and we can not get physical data $\boldsymbol{\tau}_{k}$.
    To quantify the estimation error of underlying true model parameters $\boldsymbol{\theta}^{c}$, we use the parametric bootstrapping. Based on MLE and data $\boldsymbol{\tau}_{1},\ldots,\boldsymbol{\tau}_{k-1}$, we have $\widehat{\boldsymbol{\theta}}_{k-1}$ as the estimator for $\boldsymbol{\theta}^{c}$. Then we can generate $l$ bootstrap samples, denoted as $\tilde{\boldsymbol{\theta}}^{1}_{k-1}$, $\tilde{\boldsymbol{\theta}}^{2}_{k-1},\ldots,$ $\tilde{\boldsymbol{\theta}}^{l}_{k-1}$. Then we can take the following estimation
    $f_{t,m}(\boldsymbol{\theta}^{c})\approx\frac{1}{l}\sum_{n=1}^{l}f_{t,m}(\tilde{\boldsymbol{\theta}}_{k-1}^{n}),\quad g_{t,mm}(\boldsymbol{\theta}^{c})\approx\frac{1}{l}\sum_{n=1}^{l}g_{t,mm}(\tilde{\boldsymbol{\theta}}_{k-1}^{n}).$

### v2

@@ 3 Digital Twin Calibration for Bio-SoS Mechanistic Model
To facilitate digital twin calibration, in this paper, we consider the physical system as a finite horizon stochastic process. Its dynamics can be characterized by a Bio-SoS mechanistic model specified by Eq. (1) and (2) with underlying true parameters, denoted by $\boldsymbol{\theta}^{c}=[\boldsymbol{\alpha}_{0}^{c},\boldsymbol{\alpha}_{1}^{c},\boldsymbol{\beta}^{c}]$. Our goal is to develop a calibration method that can efficiently guide the DoEs and most informative data collection to improve model fidelity and the prediction accuracy of the digital twin. For simplification, we consider batch-based cell culture experiments with the selection of initial concentration $\boldsymbol{s}_{0}$ as the decision variable.
The trajectory $\tau=(\boldsymbol{s}_{0},\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{T})$ over a horizon of $T$ time steps has the joint probability density function $p(\boldsymbol{\tau}|\boldsymbol{\theta},\boldsymbol{s}_{0})=\prod_{t=0}^{T-1}p(\boldsymbol{s}_{t+1}|\boldsymbol{s}_{t};\boldsymbol{\theta})$ with $\boldsymbol{s}_{0}=\pi(\boldsymbol{\omega})$, the calibration policy parameterized by $\boldsymbol{\omega}$. Each transition probability $p(\boldsymbol{s}_{t+1}|\boldsymbol{s}_{t};\boldsymbol{\theta})$ follows a normal distribution as defined in Eq. (4). At each $k$-th calibration iteration, we select one design following the latest policy, i.e. $\boldsymbol{s}_{0}^{(k)}=\pi(\boldsymbol{\omega}_{k})$, run an experiment to collect one sample path from the physical system and update the parameter $\widehat{\boldsymbol{\theta}}_{k}$. Specifically, given the historical samples $\mathcal{D}_{k}=\{\boldsymbol{\tau}_{n}\}_{n=1}^{k}$, we consider maximizing the log-likelihood for $\boldsymbol{\theta}$,
@@
For digital twin calibration, we select the new design of experiment at the beginning of $k$-th iteration. This selection is based on the evaluation of our current model on a set of prediction points $\mathcal{D}_{test}=\{\boldsymbol{s}_{0}^{h}\}_{h=1}^{H}$ sampled from a pre-specified distribution. The objective is to minimize the MSE of digital twin prediction,
@@
where $\boldsymbol{u}_{k}(\widehat{\boldsymbol{\theta}}_{k-1},\boldsymbol{\omega})$ denote the virtual model update function with the simulated trajectory $\boldsymbol{\tau}^{\text{d}}_{k}$ (see Eq. (8)), $Y=\phi(\boldsymbol{s}_{T})$ represents the desired system output, such as yield and product quality attributes. The superscript "p" represents for physical system and "d" for digital twin.
To sample efficiently guide digital twin calibration, we propose a gradient-based optimal learning approach with the procedure illustration as shown in Figure 2 and develop a calibration algorithm, which is summarized as Algorithm 1. In specific, for optimization problem (5), we use gradient-based method
$\widehat{\boldsymbol{\theta}}_{k}\leftarrow\widehat{\boldsymbol{\theta}}_{k-1}+\text{Adam}_{k}(\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{\tau}_{k};\widehat{\boldsymbol{\theta}}_{k-1},\boldsymbol{s}_{0}^{(k)})),$ (7)
@@
$\boldsymbol{u}_{k}(\boldsymbol{\theta},\boldsymbol{\omega})=\boldsymbol{\theta}+\text{Adam}_{k}\left(\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{\tau}_{k}^{\text{d}};\boldsymbol{\theta},\boldsymbol{\omega})\right)$ (8)
@@
Figure 2: The procedure illustration of the proposed calibration approach.
@@
Since $\boldsymbol{\theta}^{c}$ is unknown, we use the bootstrapping to quantify the model parameter estimation uncertainty in Section 5. Following the upper part of Figure 2, the prediction error propagates from $\boldsymbol{\omega}$ to model parameter update $\boldsymbol{u_{k}}$, and then to our objective $J_{k}(\boldsymbol{\omega})$. By taking the gradient of $J_{k}(\boldsymbol{\omega})$, we can discern the influence of $\boldsymbol{\omega}$ on the final prediction error, which assists in selecting a more effective design. Since deriving an exact expression for this gradient is complex, we develop an approximate closed-form representation, which will be detailed in Section 5. This approximation provides an efficient method for improving calibration policy.
@@ 4 Model Inference and Parameter Update
The gradient for the first term is $\nabla_{\boldsymbol{\theta}}\log|\Sigma(\boldsymbol{\theta})|=\text{tr}\left(\Sigma(\boldsymbol{\theta})^{-1}\nabla_{\boldsymbol{\theta}}\Sigma(\boldsymbol{\theta})\right),$ where $\nabla_{\boldsymbol{\theta}}\Sigma(\boldsymbol{\theta})$ is a tensor of $M\times M\times q$ dimension with $q$ as the length of $\boldsymbol{\theta}$, and tr()($\cdot$)( ⋅ ) represents the trace for first two dimensions. The gradient for the second term is
${\nabla_{\boldsymbol{\theta}}}(\boldsymbol{s}_{t+1}-\boldsymbol{s}_{t}-\boldsymbol{\mu}(\boldsymbol{\theta}))^{\top}\Sigma(\boldsymbol{\theta})^{-1}(\boldsymbol{s}_{t+1}-\boldsymbol{s}_{t}-\boldsymbol{\mu}(\boldsymbol{\theta}))=-2\nabla_{\boldsymbol{\theta}}{\boldsymbol{\mu}}(\boldsymbol{\theta})^{\top}\Sigma(\boldsymbol{\theta})^{-1}(\boldsymbol{s}_{t+1}-\boldsymbol{s}_{t}-\boldsymbol{\mu}(\boldsymbol{\theta}))$
$+(\boldsymbol{s}_{t+1}-\boldsymbol{s}_{t}-\boldsymbol{\mu}(\boldsymbol{\theta}))^{\top}\left(-\Sigma(\boldsymbol{\theta})^{-1}{\nabla_{\boldsymbol{\theta}}\Sigma(\boldsymbol{\theta})}\Sigma(\boldsymbol{\theta})^{-1}\right)(\boldsymbol{s}_{t+1}-\boldsymbol{s}_{t}-\boldsymbol{\mu}(\boldsymbol{\theta})).$
Given the calculated gradient, we can use the Adam optimizer (Kingma and Ba, 2014) for updating $\widehat{\boldsymbol{\theta}}$. Following the algorithm in their paper, for the $k$-th iteration, we have
@@
where $\lambda$ is the learning rate, $G_{l}=\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{\tau}_{l};\widehat{\boldsymbol{\theta}}_{l-1},\boldsymbol{s}_{0}^{(l)})$ is the gradient of the likelihood function at $l$-th iteration, $g_{l}^{2}$ is the element-wise square, $\xi_{1}=0.9,\xi_{2}=0.999$ are hyper-parameters related to exponential decay rates, and $\epsilon=10^{-8}$ is a small number to avoid zero in the denominator.
@@ 5 Policy Gradient Estimation and Optimal Learning
In this section, we turn our focus on learning the calibration policy by minimizing the prediction MSE in (6). To develop a surrogate model of this MSE, we employ the LNA in Section 5.1 to approximate the stochastic dynamics of the system, reflecting its sensitivity to initial conditions and parameter values. This approach allows us to describe the system state at any given time as a normally distributed random variable, facilitating the derivation of analytical expressions for the gradient estimator. To keep the mechanism information, in Section 5.2, a first-order Euler’s Method is utilized to derive the closed-form solutions of the SDE-based mechanisms obtained from LNA.
@@ 5.1 Linear Noise Approximation on Bio-SoS Dynamics
We use LNA (Fearnhead et al., 2014) to derive the solution of SDEs in Lemma 2, which allows us to analyze the estimation error propagation from ${\boldsymbol{\theta}}$ to the output prediction and develop a surrogate model of the MSE (6). Conditional on $\boldsymbol{s}_{0}$ and $\boldsymbol{\theta}$, we denote the state of systems at time $t$ as $\boldsymbol{s}_{t}(\boldsymbol{\theta})$. Considering the SDEs characterizing the Bio-SoS dynamics as shown in (3), we have
@@
Obtaining an analytical solution for the series of ODEs described in (13) is challenging. Therefore, we use a numerical solution approach, as described in Section 5.2. This method involves applying the solution iteratively at each time step $t$, allowing us to approximate the system’s behavior over time. For the physical system and digital twin, we denote $\boldsymbol{s}_{t}(\boldsymbol{\theta}^{c})\sim N\left(\boldsymbol{\eta}_{t},\boldsymbol{\Psi}_{t}\right)$ and $\boldsymbol{s}_{t}({\boldsymbol{u}}_{k})\sim N\left(\widehat{\boldsymbol{\eta}}_{t},\widehat{\boldsymbol{\Psi}}_{t}\right).$ Then we have $\boldsymbol{s}_{T}(\boldsymbol{\theta}^{c})-\mathbf{s}_{T}({\boldsymbol{u}}_{k})\sim N\left(\boldsymbol{\eta}_{T}-\widehat{\boldsymbol{\eta}}_{T},\boldsymbol{\Psi}_{T}+\widehat{\boldsymbol{\Psi}}_{T}\right).$ For any $m$-th component $Y$, the surrogate model of the MSE (6) is given by
${\mathbb{E}}\left[\left(Y^{\text{p}}(\boldsymbol{\theta}^{c})-Y^{\text{d}}({\boldsymbol{u}}_{k})\right)^{2}\right]=\left[\boldsymbol{\eta}_{T,m}-\widehat{\boldsymbol{\eta}}_{T,m}\right]^{2}+\boldsymbol{\Psi}_{T,mm}+\widehat{\boldsymbol{\Psi}}_{T,mm}.$ (14)
@@ 5.2 First-Order Euler’s Method for Solving LNA
Now conditional on $\widehat{\boldsymbol{\theta}}_{k-1}$, by applying a chain rule, we can update the policy parameters by
@@
where $\tilde{G}_{k}=\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{\tau}_{k}^{\text{d}};\widehat{\boldsymbol{\theta}}_{k-1},\boldsymbol{s}_{0}^{(k)})$ and $\tilde{G}_{t}=G_{t},\text{for }t=1,\ldots,k-1$. Here $\tilde{G}_{k}$ is a gradient based on simulation trajectory $\boldsymbol{\tau}_{k}^{\text{d}}$ since $\boldsymbol{\omega}_{k}$ is updated at the beginning of $k$-th iteration and we can not get physical data $\boldsymbol{\tau}_{k}$.
For the estimation of $\boldsymbol{\theta}^{c}$, we use the parametric bootstrapping. Based on MLE and data $\boldsymbol{\tau}_{1},\ldots,\boldsymbol{\tau}_{k-1}$, we have $\widehat{\boldsymbol{\theta}}_{k-1}$ as the estimator for $\boldsymbol{\theta}^{c}$. Then we can generate $l$ bootstrap samples, denoted as $\tilde{\boldsymbol{\theta}}^{1}_{k-1}$, $\tilde{\boldsymbol{\theta}}^{2}_{k-1},\ldots,$ $\tilde{\boldsymbol{\theta}}^{l}_{k-1}$. Then we can take the estimation $f_{t,m}(\boldsymbol{\theta}^{c})\approx\frac{1}{l}\sum_{n=1}^{l}f_{t,m}(\tilde{\boldsymbol{\theta}}_{k-1}^{n}),\quad g_{t,mm}(\boldsymbol{\theta}^{c})\approx\frac{1}{l}\sum_{n=1}^{l}g_{t,mm}(\tilde{\boldsymbol{\theta}}_{k-1}^{n}).$ Finally, we summarize the algorithm as follows.
For $k=1:K$
  1. 1.
  • 2.
  • Algorithm 1 Gradient-based Optimal Learning